A,B and C can finish a job in 6 days. If B and C work together,the job will take 9 days, if A anc C work alone the job will take 8 days. In how many days each man working alone do the job? Thank you.
This is what I have so far
1/x + 1/y + 1/z = 1/6
1/y + 1/z = 1/9
1/x + 1/z = 1/8
This is where I got lost
Help please thank you
: A,B and C can finish a job in 6 days. If B and C work together,the job will take 9 days, if A anc C work alone the job will take 8 days. In how many days each man working alone do the job? Thank you.
This is what I have so far
1/x + 1/y + 1/z = 1/6
1/y + 1/z = 1/9
1/x + 1/z = 1/8
This is where I got lost
Help please thank you

Answer by scott8148(2761)

you have the right equations...just substitute

substituting 2nd equation into 1st __ 1/x+1/9=1/6
__ multiplying by 18x __ 18+2x=3x __ 18=x

substituting 3rd equation into 1st __ 1/y+1/8=1/6
__ multiplying by 24y __ 24+3y=4y __ 24=y

substituting __ 1/18+1/24+1/z=1/6
__ multiplying by 72z __ 4z+3z+72=12z __ 72=5z __ 14.4=z